# chebgreen—: Learning and Interpolating Continuous Empirical Green’s Functions from Data

### v2

@@ Abstract
In this work, we present a mesh-independent, data-driven library, chebgreen, to mathematically model one-dimensional systems, possessing an associated control parameter, and whose governing partial differential equation is unknown. The proposed method learns an Empirical Green’s Function for the associated, but hidden, boundary value problem, in the form of a Rational Neural Network from which we subsequently construct a bivariate representation in a Chebyshev basis. We uncover the Green’s function, at an unseen control parameter value, by interpolating the left and right singular functions within a suitable library, expressed as points on a manifold of Quasimatrices, while the associated singular values are interpolated with Lagrange polynomials.
@@ 1 Introduction
Partial differential Equations (PDEs) have always had a pivotal role in succinctly describing natural phenomena. They allow us to mathematically express the governing fundamental laws, as rate forms [1]. Researchers have worked on data-driven approaches and deep learning techniques to solve PDEs for various systems [2, 3, 4, 5, 6, 7]. Although they hold significant practical importance, the exact PDEs governing many important phenomena remain unknown across fields, including physics, chemistry, biology, and materials science. Thus, there also has been significant work in discovering [8, 9, 10, 11, 12, 13, 14, 15, 16] the underlying PDEs from observational data. Operator Learning represents another significant research avenue; aiming to approximate the solution operator associated with a hidden partial differential equation [17]. Notable contributions in this field include, but are not limited to, Deep Operator Networks [18], Fourier Neural Operators [19], Graph Neural Operator [20], Multipole Graph Neural Operator [21], and Geometry-informed Neural Operator [22]. A detailed overview of the topic can be found in the guide on operator learning [17].
@@
In this paper, our focus is on approximating the Green’s function corresponding to some unknown governing differential equations for a given system of practical interest [23]. There has been recent theoretical work on elucidating learning rates, etc. for Green’s functions learned from data [24, 25, 26], as pertains to systems governed by elliptic or parabolic PDEs. Researchers have also utilized Deep Neural Networks to learn the Green’s function in cases where the underlying operators are weakly non-linear. This approach involves employing a dual auto-encoder architecture to uncover latent spaces where lifting the response data somewhat linearizes the problem, thereby making it more tractable for Green’s function discovery [27]. The field has also advanced through the development of Rational Neural Networks [28], which have been applied to uncover Green’s functions, offering mechanistic insights into the underlying system [29].
@@
The work on learning Empirical Green’s Functions [30] introduces a method for using observational data to learn Green’s functions, in self-adjoint contexts, without relying on machine learning techniques. The approach described in that work proposes learning Empirical Green’s function (EGF) from data in a discrete form,
@@
$\bm{G}\in\mathbb{R}^{N_{\text{sensors}}\times N_{\text{sensors}}}$ $\bm{U}\in\mathbb{R}^{N_{\text{sensors}}\times K},\bm{\Sigma}\in\mathbb{R}^{K\times K}$
That same work also proposes a methodology for interpolating these discrete EGF in a principled manner using a manifold interpolation scheme, which derives from the interpolating method for Reduced-order models [31].
In the present work, we utilize the deep learning library, greenlearning, to learn Green’s functions from data [29], and chebfun, the open-source package for computing with functions to high accuracy, in order to improve the work on interpolating Empirical Green’s Functions [30] for 1D linear operators by:
@@
With greenlearning, we learn the green’s function for a given problem using a synthetic dataset which consists of pairs of forcing functions and the system responses corresponding to those forcing functions. We learn a Neural network, $\mathcal{N}(x,y)$, which approximates the Green’s function $G(x,y)$, $\forall x,y\in\Omega\subset\mathbb{R}$ We propose creating a low-rank approximation for the neural network, $\mathcal{N}(x,y)$, using our Python implementation of chebfun2 (chebfun in two dimensions [32]), which approximates bivariate functions using Chebyshev polynomials. chebfun2 allows us to construct a Singular Value Expansion (SVE) for the learned Neural network approximating the Green’s function in the form:
@@
Note that $\mathcal{U}$ and $\mathcal{V}$ are Quasimatrices (a “matrix” in which one of the dimensions is discrete but the other is continuous), and $(\quad)^{*}$ denotes the adjoint of the Quasimatrix. Using this SVE, we interpolate the learned Green’s function using an analogue of our interpolation algorithm from the previous work [30]. This representation, combined with the proposed algorithm, enables interpolation of the Green’s function without discretizing the space, a requirement in previous work [30]. Additionally, it allows storing the Green’s function in a form suitable for high-precision computing [33]. All of this is implemented as a Python package, called chebgreen. Fig. 1 provides a brief overview of the proposed method.
@@
Figure 1: Schematic of the method for approximating Green’s functions with chebgreen—: First, we approximate the Green’s function from data using greenlearning— [29]; In the second step, our Python— implementation of chebfun— is used to compute a Singular Value Expansion for the learned Green’s function.
In Section 2, we provide a brief introduction to Green’s functions, outline the process of generating our datasets, and present the proposed methods for learning and interpolating Green’s functions. Section 3 features various numerical benchmarks for the method in one dimension, both in the presence and absence of noise. Lastly, concluding remarks are presented in Section 4.
@@ 2 Methodology
$\mathcal{L}_{\bm{\theta}}u$ $=f,$ on $\text{on }\Omega,$ (3)
$\mathcal{B}u$ $=g,$ in $\text{in }\partial\Omega,$
where, $\mathcal{B}$ is a linear differential operator specifying the boundary conditions of the problem ($g$ being the constraint on the problem boundary, $\partial\Omega$), $f$ is a forcing (or source) term, and $u$ is the unknown system response. It is pointed out that, while only a single boundary condition is shown in Eq. 3, in an actual problem setting there would be a sufficient number of such conditions to ensure well-posedness. Under suitable conditions on the operator, there exists a Green’s function $G_{\bm{\theta}}:\Omega\times\Omega\to\mathbb{R}$ associated with $\mathcal{L}_{\bm{\theta}}$ that is the impulse response of the linear differential operator and defined [23] as:
@@ 2.1 Generating the dataset
The datasets used to learn Green’s functions consist of $N_{\text{samples}}\geq 1$ forcing terms, $\{f_{j}\}_{j=1}^{N_{\text{samples}}}$, and the corresponding system’s responses, $\{u_{j}\}_{j=1}^{N_{\text{samples}}}$, satisfying Eq. 3. In the present work, our systems (3) are forced with random functions, sampled from a Gaussian process (GP), $\mathcal{GP}(0,K)$, with a covariance kernel, $K$, as motivated by recent theoretical results [24]. More specifically, the forcing terms are drawn from a GP, with mean zero and squared-exponential covariance kernel, $K_{\text{SE}}$, defined as:
@@
where, $\Omega=[a,b]\subset\mathbb{R}$ is the one-dimensional problem domain and $l>0$ is the length scale hyperparameter that sets the correlation length for this kernel. Similar to greenlearning [29], we define a normalized length scale parameter, $\sigma=l/(b-a)$, in order to remove the dependence on the length of the domain. This parameter, $\sigma$ is chosen to be larger than the spatial discretization at which we are numerically evaluating the forcing terms; to ensure that the random functions are resolved properly within the discrete representation employed. Additionally, $\sigma$ is specified to possess a suitable magnitude in order to ensure that the set of forcing terms is of full numerical rank. Note that other types of covariance kernels used in recent deep learning studies [34], such as Green’s functions related to Helmholtz equations, are likely to produce improved approximation results, as they inherently contain some information about the singular vectors of the operator, $\mathcal{L}_{\theta}$. This is reflected in the theoretical bounds for the randomized SVD with arbitrary covariance kernels, which show that one may obtain higher accuracy by incorporating knowledge of the leading singular vectors of the differential operator into the covariance kernel [25]. In the present work, we employ a generic squared-exponential kernel as one may not have prior information about the governing operator, $\mathcal{L}_{\theta}$, in real applications. In the case of problems with a periodic boundary conditions we draw the forcing functions from the periodic kernel:
@@ 2.2 Low-rank approximation of the Green’s function
where, $\sigma_{1}\geq\dots\geq\sigma_{K}\geq 0$ are the singular values, $\{\phi_{k}\}_{k=1}^{K}$ are the left singular functions, and $\{\psi_{k}\}_{k=1}^{K}$ are the right singular functions. Here, each term $\sigma_{k}\psi_{k}(y)\psi_{k}(x)$ is a rank 1 approximation which can be thought of as an “outer product” of two univariate functions.
@@ 2.3 Approximating Green’s functions with Neural Networks
In order to compute an efficient approximation for the Green’s function, $\mathcal{G}_{\theta}(x,y)$, we use the method developed by Boullé et al. [29]. The main idea is to train two Rational Neural Networks [28]: $\mathcal{N}_{\text{G}}(x,s)$ which approximates the Green’s function associated with the underlying linear differential, $\mathcal{L}_{\theta}$; and $\mathcal{N}_{\text{hom}}(x)$ which corresponds to the homogeneous solution associated with the boundary conditions, by minimizing the following loss over our training dataset $\{f_{i},u_{i}\}_{i=1}^{N_{\text{train}}}$:
[…]
@@
Incorporating this information in the loss function improves the fidelity of the computed left and right singular functions (described in Section 2.5) near the boundaries of the domain. We utilize Approximate Distance Functions [35] to enforce this by changing the loss function slightly to
[…]
where, $\alpha(x,s)$ is the Approximate Distance Function for a rectangular domain. To construct this, we start from the approximate distance function for a line segment [36] that joins $\bm{x}_{1}\equiv(x_{1},s_{1})$ and $\bm{x}_{2}\equiv(x_{2},s_{2})$, with a center $\bm{x}_{c}\equiv(\bm{x}_{1}+\bm{x}_{2})/2$ and length $L=\|\bm{x}_{1}-\bm{x}_{2}\|$:
@@
Figure 2: Example of the approximate distance function for a domain $[-4,2]\times[-4,2]$.
We then combine the approximate distance functions for the four segments, $\{\beta_{1},\beta_{2},\beta_{3},\beta_{4}\}$ with an R-equivalence solution that preserves normalization up to order $m=1$ of the distance function at all regular points, given by [37]:
@@
An example of the approximate distance function is shown in Fig. 2. In the cases where the problem does not have a Dirichlet boundary condition or the boundary conditions are unknown, we set $\alpha(x,s)\equiv 1$.
To compute the loss function described in (7), the integrals are discretized using a trapezoidal rule [38] over the measurement locations for the forcing terms $\{s_{i}\}_{i=1}^{N_{s}}$ and the system responses $\{x_{i}\}_{i=1}^{N_{x}}$. This gives us a bivariate function as a product of Rational Neural Network with a fixed function (which is defined purely in terms of the domain boundaries), $g(x,s)=\alpha(x,s)\mathcal{N}_{\text{G}}(x,s)$, which approximates the Green’s function associated with our problem. Now we would like to construct a singular value expansion (SVE) for this bivariate function. To do so, we implemented the necessary parts of the chebfun MATLAB library [33] in Python as chebpy2, using the 1D implementation, chebpy [39]. In the next two subsections, we describe the basic details of how chebfun constructs a SVE for bivariate functions using polynomial interpolants in a Chebyshev basis.

### v5

@@ Abstract
In this work, we present a mesh-independent, data-driven library, chebgreen, to mathematically model one-dimensional systems, possessing an associated control parameter, and whose governing partial differential equation is unknown. The proposed method learns an Empirical Green’s Function for the associated, but hidden, boundary value problem, in the form of a Rational Neural Network from which we subsequently construct a bivariate representation in a Chebyshev basis. We uncover the Green’s function, at an unseen control parameter value, by interpolating the left and right singular functions within a suitable library, expressed as points on a manifold of Quasimatrices, while the associated singular values are interpolated with Lagrange polynomials. This work improves upon prior work by extending the scope of applicability to non-self-adjoint operators and improves data efficiency.
@@ 1 Introduction
Partial differential equations (PDEs) have always had a pivotal role in succinctly describing natural phenomena. They allow us to mathematically express the governing fundamental laws, as rate forms [Feynman1967_Physical]. Researchers have worked on data-driven approaches and deep learning techniques to solve PDEs for various systems [Raissi2018_Deep, Raissi2018_Hidden, Berg2017_Neural, Lu2021_DeepXDE, Sirignano2018_DGM, Weinan2018_DeepRitz]. Although they hold significant practical importance, the exact PDEs governing many important phenomena remain unknown across fields, including physics, chemistry, biology, and materials science. Thus, there also has been significant work in discovering [Raissi2019_Physics, Schaeffer2017_learning, Brunton2016_Discovering, Champion2019_Discovery, Rudy2017_Data, Bonneville2022_Bayesian, Stephany2022_PDE_READ, Stephany2024_PDE_LEARN, Stephany2024_Weak_PDE_LEARN] the underlying PDEs from observational data. Operator Learning represents another significant research avenue; aiming to approximate the solution operator associated with a hidden partial differential equation [Boulle2024_Operator]. Notable contributions in this field include, but are not limited to, Deep Operator Networks [Lu2021_DeepONet] which learns a branch (input encoding) and trunk (output encoding) networks to efficiently and accurately learn nonlinear operators from data, Fourier Neural Operators [Li2020_FNO] which efficiently learn operators by parameterizing the integral kernel directly in Fourier space, Graph Neural Operator [Li2020_GraphNO] which focus on capturing the Green kernel’s short-range interactions to reduce the computational complexity of integration by using message passing on Graph Neural Networks, Multipole Graph Neural Operator [Li2020_MGNO] which enhance graph-based methods by decomposing the Green kernel into a sum of low-rank kernels to capture long-range dependencies, and Geometry-informed Neural Operator [Li2023_GINO], which incorporate geometric information into the operator learning framework. A detailed overview of operator learning can be found in this survey article [Boulle2024_Operator].
@@
In this paper, our focus is on approximating the Green’s function corresponding to some unknown governing differential equations for a given system of practical interest [Evans2010_PDE]. The Hilbert-Schmidt Theorem [Khanfer2024] provide a powerful framework for analyzing how Green’s functions associated with hidden differential operators, may be approximated. Within this setting, the Singular Value Expansion provides a valuable method for studying this approximation, particularly in the context of solving inverse problems and learning from data [Fasshauer2015_Kernel]. Our proposed approach to operator learning is of practical value, as it allows for the construction of data-driven models that inherently respect underlying physical laws governing a system of interest; effectively solving differential equations in a mesh-free manner. There has been recent theoretical work on elucidating learning rates, etc. for Green’s functions learned from data [Boulle2021_Theory, Boulle2021_SVD, Boulle2022_ParabolicPDE], as pertains to systems governed by elliptic or parabolic PDEs. Researchers have also utilized Deep Neural Networks to learn the Green’s function in cases where the underlying operators are weakly non-linear. This approach involves employing a dual auto-encoder architecture to uncover latent spaces where lifting the response data somewhat linearizes the problem, thereby making it more tractable for Green’s function discovery [Gin2021_DeepGreen]. The field has also advanced through the development of Rational Neural Networks [Boulle2020_Rational], which have been applied to uncover Green’s functions, offering mechanistic insights into the underlying system [Boulle2021_Greenlearning].
@@
The work on learning Empirical Green’s Functions [Praveen2023_EGF] introduces a method for using observational data to learn Green’s functions, in self-adjoint contexts, without relying on machine learning techniques. The approach described in that work proposes learning Empirical Green’s function (EGF) from data in a discrete form,
@@
$\bm{G}\in\mathbb{R}^{N_{\text{sensors}}\times N_{\text{sensors}}}$ $\text{, }\bm{U}\in\mathbb{R}^{N_{\text{sensors}}\times K},\bm{\Sigma}\in\mathbb{R}^{K\times K}$
That same work also proposes a methodology for interpolating these discrete EGF in a principled manner using a manifold interpolation scheme, which derives from the interpolating method for Reduced-order models [Amsallem2008_ROM].
In the present work, we utilize the deep learning library, greenlearning, to learn Green’s functions from data [Boulle2021_Greenlearning], and chebfun, the open-source package for computing with functions to high accuracy, in order to improve the work on interpolating Empirical Green’s Functions [Praveen2023_EGF] for 1D linear operators by:
@@
With greenlearning, we learn the green’s function for a given problem using a synthetic dataset which consists of pairs of forcing functions and the system responses corresponding to those forcing functions. We learn a Neural network, $\mathcal{N}(x,y)$, which approximates the Green’s function $G(x,y)$, $\forall x,y\in\Omega\subset\mathbb{R}$ . We propose creating a low-rank approximation for the neural network, $\mathcal{N}(x,y)$, using our Python implementation of chebfun2 (chebfun in two dimensions [Townsend_Chebfun2]), which approximates bivariate functions using Chebyshev polynomials. chebfun2 allows us to construct a Singular Value Expansion (SVE) for the learned Neural network approximating the Green’s function in the form:
@@
Note that $\mathcal{U}$ and $\mathcal{V}$ are Quasimatrices (a “matrix” in which one of the dimensions is discrete but the other is continuous), and $(\quad)^{*}$ denotes the adjoint of the Quasimatrix. Using this SVE, we interpolate the learned Green’s function using an analogue of our interpolation algorithm from the previous work [Praveen2023_EGF]. This representation, combined with the proposed algorithm, enables interpolation of the Green’s function without discretizing the space, a requirement in previous work [Praveen2023_EGF]. Additionally, it allows storing the Green’s function in a form suitable for high-precision computing [Chebfun]. All of this is implemented as a Python package, called chebgreen. Fig. 1 provides a brief overview of the proposed method.
@@
Figure 1: Schematic of the method for approximating Green’s functions with chebgreen—: First, we approximate the Green’s function from data using greenlearning— [Boulle2021_Greenlearning]; In the second step, our Python— implementation of chebfun— is used to compute a Singular Value Expansion for the learned Green’s function.
In Section 2, we provide a brief introduction to Green’s functions, outline the process of generating our datasets, and present the proposed methods for learning and interpolating Green’s functions. Section 4 features various numerical benchmarks for the method in one dimension, both in the presence and absence of noise. Lastly, concluding remarks are presented in LABEL:sec:conclusion.
@@ 2 Methodology
$\mathcal{L}_{\bm{\theta}}u$ $=f,$ $\text{on }\Omega,$ (3)
$\mathcal{B}u$ $=g,$ $\text{in }\partial\Omega,$
where, $\mathcal{B}$ is a linear differential operator specifying the boundary conditions of the problem ($g$ being the constraint on the problem boundary, $\partial\Omega$), $f$ is a forcing (or source) term, and $u$ is the unknown system response. It is pointed out that, while only a single boundary condition is shown in Eq. 3, in an actual problem setting there would be a sufficient number of such conditions to ensure well-posedness. Under suitable conditions on the operator, there exists a Green’s function $G_{\bm{\theta}}:\Omega\times\Omega\to\mathbb{R}$ associated with $\mathcal{L}_{\bm{\theta}}$ that is the impulse response of the linear differential operator and defined [Evans2010_PDE] as:
@@ 2.1 Generating the dataset
The datasets used to learn Green’s functions consist of $N_{\text{samples}}\geq 1$ forcing terms, $\{f_{j}\}_{j=1}^{N_{\text{samples}}}$, and the corresponding system’s responses, $\{u_{j}\}_{j=1}^{N_{\text{samples}}}$, satisfying Eq. 3. In the present work, our systems (3) are forced with random functions, sampled from a Gaussian process (GP), $\mathcal{GP}(0,K)$, with a covariance kernel, $K$, as motivated by recent theoretical results [Boulle2021_Theory]. More specifically, the forcing terms are drawn from a GP, with mean zero and squared-exponential covariance kernel, $K_{\text{SE}}$, defined as:
@@
where, $\Omega=[a,b]\subset\mathbb{R}$ is the one-dimensional problem domain and $l>0$ is the length scale hyperparameter that sets the correlation length for this kernel. Similar to greenlearning [Boulle2021_Greenlearning], we define a normalized length scale parameter, $\sigma=l/(b-a)$, in order to remove the dependence on the length of the domain. This parameter, $\sigma$ is chosen to be larger than the spatial discretization at which we are numerically evaluating the forcing terms; to ensure that the random functions are resolved properly within the discrete representation employed. Additionally, $\sigma$ is specified to possess a suitable magnitude in order to ensure that the set of forcing terms is of full numerical rank. Note that other types of covariance kernels used in recent deep learning studies [Li2021_Fourier], such as Green’s functions related to Helmholtz equations, are likely to produce improved approximation results, as they inherently contain some information about the singular vectors of the operator, $\mathcal{L}_{\theta}$. This is reflected in the theoretical bounds for the randomized SVD with arbitrary covariance kernels, which show that one may obtain higher accuracy by incorporating knowledge of the leading singular vectors of the differential operator into the covariance kernel [Boulle2021_SVD]. In the present work, we employ a generic squared-exponential kernel as one may not have prior information about the governing operator, $\mathcal{L}_{\theta}$, in real applications. In the case of problems with a periodic boundary conditions we draw the forcing functions from the periodic kernel:
@@ 2.2 Low-rank approximation of the Green’s function
where, $\sigma_{1}\geq\dots\geq\sigma_{K}\geq 0$ are the singular values, $\{\phi_{k}\}_{k=1}^{K}$ are the left singular functions, and $\{\psi_{k}\}_{k=1}^{K}$ are the right singular functions. Here, each term $\sigma_{k}\phi_{k}(y)\psi_{k}(x)$ is a rank 1 approximation which can be thought of as an “outer product” of two univariate functions.
We choose a truncated Singular Value Expansion (SVE) because the SVE of an infinite-dimensional Hilbert-Schmidt (HS) operator so long as the Green’s function is square-integrable), after retaining $K$ terms, yields the best rank-$K$ approximation under the HS and operator norms; by the Eckart-Young-Mirsky theorem [Stewart1990_SVE]. This representation is important because it allows us to interpolate learned Green’s functions between different parametric values (Section 2.7). Moreover, in practice, we learn a rational approximation to the Green’s function as described in the next section, which has exponentially decaying singular values. This makes an SVE the appropriate choice for representing a learned Green’s functions.
@@ 2.3 Approximating Green’s functions with Neural Networks
In order to compute an efficient approximation for the Green’s function, $\mathcal{G}_{\theta}(x,y)$, we use the method developed by Boullé et al. [Boulle2021_Greenlearning]. The main idea is to train two Rational Neural Networks [Boulle2020_Rational]: $\mathcal{N}_{\text{G}}(x,s)$ which approximates the Green’s function associated with the underlying linear differential, $\mathcal{L}_{\theta}$; and $\mathcal{N}_{\text{hom}}(x)$ which corresponds to the homogeneous solution associated with the boundary conditions, by minimizing the following loss over our training dataset $\{f_{i},u_{i}\}_{i=1}^{N_{\text{train}}}$:
$\text{Loss}=\frac{1}{N_{\text{train}}}\sum_{i=1}^{N_{\text{train}}}\frac{1}{\|u_{i}\|_{L^{2}(\Omega)}^{2}}\int_{\Omega}\left(u_{i}(x)-\mathcal{N}_{\text{hom}}(x)-\int_{\Omega}\mathcal{N}_{\text{G}}(x,s)f_{i}(s)ds\right)^{2}dx$ (6)
Rational Neural Networks offer a significant advantage for learning Green’s functions: their poles offer a natural means for approximating singularities within underlying Green’s functions and corresponding homogeneous solutions [Boulle2021_Greenlearning]. Additionally, their superior approximation properties, validated both theoretically and empirically [Boulle2020_Rational], make them particularly well-suited for this task.
@@
Incorporating this information in the loss function improves the fidelity of the computed left and right singular functions (described in Section 2.5) near the boundaries of the domain. We utilize Approximate Distance Functions [Sukumar2022_Distancefunctions] to enforce this by changing the loss function slightly to
$\text{Loss}=\frac{1}{N_{\text{train}}}\sum_{i=1}^{N_{\text{train}}}\frac{1}{\|u_{i}\|_{L^{2}(\Omega)}^{2}}\int_{\Omega}\left(u_{i}(x)-\mathcal{N}_{\text{hom}}(x)-\int_{\Omega}\alpha(x,s)\mathcal{N}_{\text{G}}(x,s)f_{i}(s)ds\right)^{2}dx,$ (7)
where, $\alpha(x,s)$ is the Approximate Distance Function for a rectangular domain. To construct this, we start from the approximate distance function for a line segment [Rvachev2001_LineADF] that joins $\bm{x}_{1}\equiv(x_{1},s_{1})$ and $\bm{x}_{2}\equiv(x_{2},s_{2})$, with a center $\bm{x}_{c}\equiv(\bm{x}_{1}+\bm{x}_{2})/2$ and length $L=\|\bm{x}_{1}-\bm{x}_{2}\|$:
@@
Figure 2: Example of the approximate distance function for a domain $4,2]\times[-4,2]$.
We then combine the approximate distance functions for the four segments, $\{\beta_{1},\beta_{2},\beta_{3},\beta_{4}\}$ with an R-equivalence solution that preserves normalization up to order $m=1$ of the distance function at all regular points, given by [Biswas2004_Requiv]:
@@
An example of the approximate distance function is shown in Fig. 2. In the cases where the problem does not have a Dirichlet boundary condition or the boundary conditions are unknown, we set $\alpha(x,s)\equiv 1$. In this context, “approximate” refers to distance functions approximating the exact signed distance function, which are, themselves, the orthogonal distance from a point to the boundary of a set. The adopted approximation implicitly specifies a Dirichlet boundary condition. However, this is not a requirement; Rational Neural Networks can learn the boundary conditions directly, from training data [Boulle2021_Greenlearning]. When Dirichlet boundary conditions are known a priori, employing approximate distance functions improves the numerical approximation of the Green’s functions near the boundaries, which in turn improves the numerical approximation of the learned singular functions (Section 2.5).
To compute the loss function described in (7), the integrals are discretized using a trapezoidal rule [Endre2003_NumericalAnalysis] over the measurement locations for the forcing terms $\{s_{i}\}_{i=1}^{N_{s}}$ and the system responses $\{x_{i}\}_{i=1}^{N_{x}}$. The trapezoidal rule is used for numerical integration because the numerical experiments are done on a uniform grid. However, a uniform grid is not a requirement; the underlying implementation allows users to specify arbitrary sensor locations for both forcing terms and system responses as detailed in Section 2.1. Alternatively, users can leverage the implementation Monte Carlo integration in the library if sensor locations are non-uniformly distributed. Prior work on learning Green’s functions with Rational Neural Networks [Boulle2021_Greenlearning] has demonstrated the robustness of Rational Neural Networks to variations in sensor placement and the choice of quadrature rule for integration. This gives us a bivariate function as a product of Rational Neural Network with a fixed function (which is defined purely in terms of the domain boundaries), $g(x,s)=\alpha(x,s)\mathcal{N}_{\text{G}}(x,s)$, which approximates the Green’s function associated with our problem. Now we would like to construct a singular value expansion (SVE) for this bivariate function. To do so, we implemented the necessary parts of the chebfun MATLAB library [Chebfun] in Python as chebpy2, using the 1D implementation, chebpy [Chebpy]. In the next two subsections, we describe the basic details of how chebfun constructs a SVE for bivariate functions using polynomial interpolants in a Chebyshev basis.